OpenStudy (anonymous):

(7,-7),m=-3 written in slope intercept form

OpenStudy (anonymous):

hi. have you come to any epiphanies? or could you still use some help?

OpenStudy (anonymous):

I need help please! :)

OpenStudy (anonymous):

okeedokee. Any clues as to how to write this little puppy?

OpenStudy (anonymous):

lets start with y-y1=m(x-x1)

OpenStudy (anonymous):

alright...

OpenStudy (anonymous):

now they give you a point with coordinates (7,-7) and m=-3. So now, you substitute the coordinates and the slope into the above equation

OpenStudy (anonymous):

does that help or are you still confused?

OpenStudy (anonymous):

well, substitution leads to y-(-7)=-3(x-7)

OpenStudy (anonymous):

i think you can do the rest... take care

OpenStudy (anonymous):

whenever you are given the coordinates of a point and the slope of a line, you can use the above form to find the equation of the line.

OpenStudy (anonymous):

you can also use the slope intercept form, y=mx+b, but it would take longer. you would have to solve for b.